### v1

# Mass-imbalance effect on the cluster formation in a one-dimensional Fermi gas with coexistent $s$- and $p$-wave interactions

Abstract
We consider the mass-imbalance effect on the clustering in a one-dimensional two-component Fermi gas with coexistent even- and odd-wave interactions resulting in different configurations of clustering phases. We obtain the solutions of both stable two- and three-body cluster states with different mass ratios and configurations by solving the corresponding variational equations. We feature out phase diagrams consisting of the $s$- and $p$-wave pairing phases, and tripling phase with different configurations, in a plane of $s$- and $p$-wave pairing strengths. As for the in-vacuum case, the three-body clustering is always the lowest-lying phase. While for the in-medium case, the Cooper tripling phase dominates over the pairing phases when both $s$- and $p$-wave interactions are moderately strong. There is also a competition between different clustering configurations of three-body clustering.
I Introduction
The investigation of superconductors and superfluids has been one of the frontiers in modern physics. It is indicated that the competition and the coexistence of various orders are worth being investigated [1, 2, 3, 4, 5], which have important applications in different fields such as condensed-matter and nuclear physics. However, to understand the microscopic mechanism of non-trivial states arising in the mass-imbalanced systems with coexistent orders is still one of the most challenging open questions.
It would be useful to study the mass-imbalance effect on multibody clustering in various systems. The Efimov state [6] has attracted lots of interest in ultracold atoms, nuclear physics, and so on. In the corresponding investigation, experimental evidence of Efimov physics has been reported in different mass-imbalanced ultracold atomic gas mixtures such as 41K-87Rb [7] or 6Li-133Cs [8] ones. The two-nucleon halo nuclei such as 11Li (9Li core ++ 22 halo neutrons) is also regarded as an ideal candidate to experimentally detect the Efimov states [9]. So far, various possible superconductors with competing odd- and even-parity pairing channels have been experimentally realized in the mass-imbalanced systems, including UTe2 [4, 10], UPt3 [11], WTe2 [13], and Sr2RuO4 [14, 15]. Moreover, in nuclear physics, there is also a novel state called hypernucleus consisting of at least one hyperon in addition to the normal nucleons [16]. It shows many interesting properties, such as that hypernuclei containing the lightest hyperon $\Lambda$ (such as $^{7}_{\Lambda}$Li) tend to be more tightly bound than normal nuclei [17]. The hypernuclei with small particle number would be useful for the preliminary studies.
The ultracold atomic gas is regarded as an ideal platform for investigating such non-trivial quantum many-body phenomena, due to its tunnable interactions and densities [18, 19]. Both $s$- and $p$-wave interactions are tunable near the Feshbach resonance [20, 21, 22, 23, 19, 24, 25]. Consequently, it is interesting to realize the situation that Fermi superfluids with hybridized $s$- and $p$-wave pairings in cold atomic systems via overlapped Feshbach resonances [18, 26, 27]. In such a system, in addition to $s$- and $p$-wave Cooper pairings, the clustering associated with the Cooper instability may be extended to induce the bound states consisting of three particles or even more. In order to study such a many-body problem, the generalized Cooper problem has been further applied to cluster states such as three-body Cooper triples [28, 29, 30, 31, 32, 5, 33] and even four-body Cooper quartets [34, 35, 36, 37, 38, 39]. On the other hand, with the help of specific geometry and trap potential, low dimensions and lattice structures can be achieved in cold-atomic systems [40, 41, 42, 43]. Such low-dimensional systems in condensed-matter or nuclear physics have also attracted much attention [44, 45, 46].
In the previous work [5], we studied the degenerate two-component fermions with coexistent interspecies and intraspecies interactions, which lead to the Cooper instabilities towards the $s$- and $p$-wave Cooper pairs, and the Cooper triple with mass-balanced configuration. However, the mass imbalance is also worth being investigated, and the intraspecies interaction is not necessary limited to exist only within one component. In this work, we are going to figure out the properties in a system with both $s$- and $p$-wave interactions as shown in Fig. 1. The two- and three-body clusterings with different configurations are also anticipated. In Ref. [47], by activating the orbitals degree of freedom, the authors reported that the emergence of even-parity ($s$-wave) interaction in a spin-polarized fermionic 40K gas besides the odd-wave ($p$-wave) interactions among identical fermions in each orbital. Based on that, present model setup might be achieved experimentally via fulfilling multiorbital structure with different components of atoms in the future.
Figure 1: Schematic figure representing our model. We consider the degenerate two-component mass-imbalanced fermions (states $a$ and $b$ occupying the levels with momentum $k$ up to the Fermi momentum $k_{\rm F}$) and the consequence of coexistent interspecies $s$-wave interaction $V_{s}$ and intraspecies $p$-wave interactions $V_{i}$ (acting on two identical fermions in the $i=a,b$ state), which lead to the Cooper instabilities towards the $s$- and $p$-wave Cooper pairs, and the Cooper triples.
In this paper, we investigate the fate of in-vacuum and in-medium two- and three-body correlations in a one-dimensional two-component Fermi gas with coexistent $s$- and $p$-wave interactions. Specific attention is paid to the mass imbalance and different configurations of clustering. In general, the mass of two components of fermions are adopted as different ones, and $p$-wave interactions among both kinds of atoms are taken into considerations. By solving the variational equations, we show the solutions of the stable clusters. We also figure out a phase diagram consisting of pairing and tripling states with different configurations.
This paper is organized as follows. The theoretical framework is presented in Sec. II, where we show the Hamiltonian for present system. We apply a variational approach for in-medium two- and three-body states on top of the Fermi sea to this model. In Sec. III.1, we first discuss our numerical results for the in-vacuum bound states and the ground-state phase diagram. The results of in-medium case are shown in Sec. III.2. Finally, a summary and perspectives will be given in Sec. IV. In the following, we take $\hbar=c=k_{\rm B}=1$. The system size is taken to be a unit.
II Theoretical framework
In this work, we consider a one-dimensional mass-imbalanced two-component fermions with coexistent short-range interactions. The intraspecies odd-parity ($p$-wave) interactions among each kind of fermions are both taken into considerations. While as for the interspecies interaction, we assume the even-parity ($s$-wave) one is dominant and neglect the odd-parity ($p$-wave) interaction one as the same as the previous work [5]. Consequently, the present system can be described by the following Hamiltonian as
$H=\,$ $K+V_{s}+V_{a}+V_{b},$ (1)
$K=\,$ $\sum_{k}\left(\xi_{k,a}a_{k}^{\dagger}a_{k}+\xi_{k,b}b_{k}^{\dagger}b_{k}\right),$ (2)
$V_{a}=\,$ $\frac{U_{a}}{2}\sum_{p,p^{\prime},q}pp^{\prime}a_{p+q/2}^{\dagger}a_{-p+q/2}^{\dagger}a_{-p^{\prime}+q/2}a_{p^{\prime}+q/2},$ (3)
$V_{b}=\,$ $\frac{U_{b}}{2}\sum_{p,p^{\prime},q}pp^{\prime}b_{p+q/2}^{\dagger}b_{-p+q/2}^{\dagger}b_{-p^{\prime}+q/2}b_{p^{\prime}+q/2},$ (4)
$V_{s}=\,$ $U_{s}\sum_{p,p^{\prime},q}a_{p+\frac{m_{r}}{m_{b}}q}^{\dagger}b_{-p+\frac{m_{r}}{m_{a}}q}^{\dagger}b_{-p^{\prime}+\frac{m_{r}}{m_{a}}q}a_{p^{\prime}+\frac{m_{r}}{m_{b}}q},$ (5)
where $\xi_{k,i}=k^{2}/(2m_{i})-\mu_{i}$ ($i=a,b$) in the kinetic term $K$ is the single-particle energy with a momentum $k$, an atomic mass $m_{i}$, and chemical potential $\mu_{i}$. Here the reduced mass reads $m_{r}=\frac{m_{a}m_{b}}{m_{a}+m_{b}}$. Note that in general the cases with $m_{a}\neq m_{b}$ will be considered in this paper. In the generalized Cooper problems, we take $\mu_{i}=E_{\textrm{F},i}$ where $E_{\textrm{F},i}$ is the Fermi energy of species $i$. Due to the mass difference, there is a mismatch of the Fermi energy (or chemical potential) between two different kinds of fermions, namely, $\mu_{a}\neq\mu_{b}$. $V_{i}$ represents the two-body $p$-wave interaction between two identical atoms $i$ with a coupling constant $U_{i}$, and $V_{s}$ corresponds to the two-body $s$-wave interaction with a coupling constant $U_{s}$. The contact couplings $U_{s}$ and $U_{i}$ can be renormalized by introducing the $s$-wave and $p$-wave scattering lengths as
$U_{s}=-\frac{1}{m_{r}a_{s}},$ (6)
and
$\frac{m_{i}}{2a_{p}}=\frac{1}{U_{i}}+\sum_{p}\frac{p^{2}}{2\varepsilon_{p,i}}$ (7)
with $\varepsilon_{p,i}=p^{2}/(2m_{i})$, respectively. Here we assume the $p$-wave scattering lengths of two components are always tuned to be equal to each other. Incidentally, we are interested in an attractive $s$-wave interaction in this work, the positive $s$-wave scattering length $a_{s}>0$ is taken.
As schematically shown in Fig. 1, with the mass imbalance and the $p$-wave interactions among both kinds of atoms, there are in total five phases of correlations expected to appear in the present system. In detail, as for the pairing states, we have $s$-wave pairing, $p$-wave pairing ($aa$ configuration) and $p$-wave pairing ($bb$ configuration). While for the tripling states, we have aabaab and abbabb configurations, respectively.
The trial wave functions for these in-medium two-and three-body states are adopted as
$|\Psi_{ab}\rangle=\,$ $\sum_{k}\theta(|k|-k_{\rm F})\Omega^{ab}_{k}a_{k}^{\dagger}b_{-k}^{\dagger}|{\rm FS}\rangle$ (8)
$=\,$ $\sum_{k}^{\prime}\Omega^{ab}_{k}a_{k}^{\dagger}b_{-k}^{\dagger}|{\rm FS}\rangle,$ (9)
for $s$-wave pairing state [48],
$|\Psi_{aa}\rangle=\,$ $\sum_{k}\theta(|k|-k_{\rm F})\Omega^{aa}_{k}a_{k}^{\dagger}a_{-k}^{\dagger}|{\rm FS}\rangle$ (10)
$=\,$ $\sum_{k}^{\prime}\Omega^{aa}_{k}a_{k}^{\dagger}a_{-k}^{\dagger}|{\rm FS}\rangle,$ (11)
for $p$-wave pairing ($aa$ configuration) state [32],
$|\Psi_{bb}\rangle=\,$ $\sum_{k}\theta(|k|-k_{\rm F})\Omega^{bb}_{k}b_{k}^{\dagger}b_{-k}^{\dagger}|{\rm FS}\rangle$ (12)
$=\,$ $\sum_{k}^{\prime}\Omega^{bb}_{k}b_{k}^{\dagger}b_{-k}^{\dagger}|{\rm FS}\rangle,$ (13)
for $p$-wave pairing ($bb$ configuration) state,
$|\Psi_{aab}\rangle=\,$ $\sum_{p,q}\theta(|p+q/2|-k_{\rm F})\theta(|-p+q/2|-k_{\rm F})$
$\times\theta(|-q|-k_{\rm F})\Omega^{aab}_{p,q}a_{p+q/2}^{\dagger}a_{-p+q/2}^{\dagger}b_{-q}^{\dagger}|{\rm FS}\rangle$ (14)
$=\,$ $\sum_{p,q}^{\prime}\Omega^{aab}_{p,q}a_{p+q/2}^{\dagger}a_{-p+q/2}^{\dagger}b_{-q}^{\dagger}|{\rm FS}\rangle,$ (15)
for tripling (aabaab configuration) state [5], and
$|\Psi_{abb}\rangle=\,$ $\sum_{p,q}\theta(|p+q/2|-k_{\rm F})\theta(|-p+q/2|-k_{\rm F})$
$\times\theta(|-q|-k_{\rm F})\Omega^{abb}_{p,q}a_{-q}^{\dagger}b_{p+q/2}^{\dagger}b_{-p+q/2}^{\dagger}|{\rm FS}\rangle$ (16)
$=\,$ $\sum_{p,q}^{\prime}\Omega^{abb}_{p,q}a_{-q}^{\dagger}b_{p+q/2}^{\dagger}b_{-p+q/2}^{\dagger}|{\rm FS}\rangle,$ (17)
for tripling (abbabb configuration) state, respectively. In the above trial wave functions, $\Omega^{mn\cdots}_{q_{1},q_{2},\cdots}$ denotes the variational parameter for the state with $mn\cdots$ configuration and a set of momenta $q_{1},q_{2},\cdots$, and $|{\rm FS}\rangle$ represents the Fermi sea. From the parity and symmetry, it is easy to find $\Omega^{iij}_{p,q}=-\Omega^{iij}_{-p,q}$ and $\Omega_{p,q}^{iij}=\Omega_{p,-q}^{iij}$ ($i,j=a,b$ and $i\neq j$) for Cooper triple phases. In addition, here $\sum_{p_{1},p_{2},\cdots}^{\prime}$ is adopted to denote the the momentum summation restricted by the Fermi surface for convenience. The step functions associated with the Fermi surface will be recovered when we numerically evaluate the momentum summation. By minimizing the in-medium two- and three-body ground-state energies based on the variational principle, the variational parameter $\Omega^{mn\cdots}_{q_{1},q_{2},\cdots}$ will be determined correspondingly.
Calculating the expectation values of Hamiltonian associated with each trial wave function $\langle\Psi_{mn\cdots}|H|\Psi_{mn\cdots}\rangle$, and then applying the variational principle, we arrive that
$\frac{\delta}{\delta{\Omega^{mn\cdots}_{q_{1},{q_{2}},\cdots}}^{*}}\langle\Psi_{mn\cdots}|(H-E^{mn\cdots}_{\alpha})|\Psi_{mn\cdots}\rangle=0,$ (18)
where $E_{\alpha}^{mn\cdots}$ ($\alpha=2,3$) is the ground-state energy of a Cooper cluster with $mn\cdots$ configuration. As a result,the in-medium variational equations read
$1+U_{s}\sum_{q}^{\prime}\frac{1}{\xi_{q,a}+\xi_{q,b}-E_{2,s}}=0,$ (19)
for the $s$-wave pairing [48],
$1+U_{i}\sum_{p}^{\prime}\frac{p^{2}}{\xi_{p,i}+\xi_{-p,i}-E_{2,p}^{ii}}=0,$ (20)
for $p$-wave pairing ($ii$ configuration) [32],
$(\xi_{p+q/2,a}+\xi_{-p+q/2,a}+\xi_{-q,b}-E_{3}^{aab})\Omega^{aab}_{p,q}+p\Gamma^{aab}_{p}(q)$
$+\Gamma^{aab}_{s}(p-q/2)-\Gamma^{aab}_{s}(-p-q/2)=0,$ (21)
for tripling (aabaab configuration) [5], and
$(\xi_{-q,a}+\xi_{p+q/2,b}+\xi_{-p+q/2,b}-E_{3}^{abb})\Omega^{abb}_{p,q}+p\Gamma^{abb}_{p}(q)$
$+\Gamma^{abb}_{s}(p-q/2)-\Gamma^{abb}_{s}(-p-q/2)=0,$ (22)
for tripling (abbabb configuration), respectively. To simplify the ensuing expressions, in the above variational equations we introduce
$\Gamma_{p}^{iij}(q)=U_{i}\sum_{p^{\prime}}^{\prime}p^{\prime}\Omega^{iij}_{p^{\prime},q},\quad\Gamma^{iij}_{s}(k)=U_{s}\sum_{q^{\prime}}^{\prime}\Omega^{iij}_{k+q^{\prime}/2,q^{\prime}},$ (23)
with $i,j=a,b$ and $i\neq j$. We also figure out that
$\Gamma^{iij}_{p}(q)=\Gamma^{iij}_{p}(-q),\quad\Gamma^{iij}_{s}(k)=-\Gamma^{iij}_{s}(-k).$ (24)
Correspondingly, the in-medium three-body equations for $\Gamma^{iij}_{p}(q)$ and $\Gamma^{iij}_{s}(k)$ reads
$\Gamma^{iij}_{p}(q)\left[1+U_{i}\sum_{p}^{\prime}\frac{p^{2}}{\xi_{p+q/2,i}+\xi_{-p+q/2,i}+\xi_{-q,j}-E_{3}^{iij}}\right]$ (25)
$=\,$ $-U_{i}\sum_{p}^{\prime}p\frac{\Gamma^{iij}_{s}(p-q/2)-\Gamma^{iij}_{s}(-p-q/2)}{\xi_{p+q/2,i}+\xi_{-p+q/2,i}+\xi_{-q,j}-E_{3}^{iij}},$ (26)
and
$\Gamma^{iij}_{s}(k)\left[1+U_{s}\sum_{q}^{\prime}\frac{1}{\xi_{k+q,i}+\xi_{-k,i}+\xi_{-q,j}-E_{3}^{iij}}\right]$ (27)
$=\,$ $-U_{s}\sum_{q}^{\prime}\frac{(k+q/2)\Gamma^{iij}_{p}(q)-\Gamma^{iij}_{s}(-k-q)}{\xi_{k+q,i}+\xi_{-k,i}+\xi_{-q,j}-E_{3}^{iij}},$ (28)
respectively.
Before discussing the numerical demonstrations in detail, we note that present variational approach is able to describe the BCS-BEC crossover and its three-body counterpart qualitatively [49]. The variational wave functions for calculating the energies of Cooper clusters can well depict both the Cooper clustering in the weak-coupling BCS limit and the tightly bound state formation in the strong-coupling BEC limit in a unified manner.
III Results and discussion
III.1 In-vaccum case
In the absence of the Fermi sea, namely, by taking the reference state as the vacuum $|0\rangle$ instead of the Fermi sea $|\textrm{FS}\rangle$, we first investigate the in-vacuum two- and three-body clustering in present mass-imbalanced two-component Fermi gas.
Figure 2: (a). In-vacuum and in-medium $s$-wave pairing energies as functions of $s$-wave scattering length. (b). In-vacuum and in-medium $p$-wave pairing energies with different configurations as functions of $p$-wave scattering length. (c). In-vacuum and in-medium three-body energies with different configurations as functions of $p$-wave scattering length at fixed $s$-wave interaction strength as $1/(\Lambda a_{s})\times 10=2.7$ with $\Lambda$ the momentum cutoff. In all panels, mass ratio is fixed as $m_{a}/m_{b}=2$, and the reference energy scale is $E_{\Lambda}=\Lambda^{2}/(2m_{r})$.
In Fig. 2, by taking the mass ratio $m_{a}=2m_{b}$ as a typical example, we plot in-vacuum clustering energies as functions of scattering lengths. As for the in-vacuum two-body energies shown in Figs. 2(a) and (b), the ss($p$)-wave pairing one monotonically decreases with the increase of ss($p$)-wave interaction strengths. From Fig. 2(b), it is seen that the in-vacuum two-body pairing with $aa$ configuration is always suppressed by the one with $bb$ configuration. It is due to the mass difference, where the pairing consisting of lighter atoms is more tightly bound at a fixed interaction strength. Such a tendency can be qualified from Eq. (20). One has
$\frac{E_{2,p}^{ii}}{E_{\Lambda}}\propto\frac{m_{r}}{m_{i}}$ (29)
in the reference scale of $E_{\Lambda}=\Lambda^{2}/(2m_{r})$, where $\Lambda$ denotes the momentum cutoff adopted in the practical calculations. It naturally indicates $p$-wave bbbb pairing is more tightly bound than the $aa$ configuration.
[…]
Figure 3: In-vacuum phase diagram in the plane of $1/(\Lambda a_{s})$ and $1/(\Lambda a_{p})$. The mass ratio is taken as $m_{a}=2m_{b}$.
We feature out the in-vacuum phase diagram in the plane of $1/(\Lambda a_{s})$ and $1/(\Lambda a_{p})$ as shown in Fig. 3. The in-vacuum phase diagram gives the lowest-lying phase among all the five configurations of two- or three-body clusters discussed in this work. The mass ratio is taken as $m_{a}=2m_{b}$. It is seen that in the interaction regime concerned in this paper, $|E_{3}|>|E_{2}|$ always holds regardless of specific two- and three-body configurations in the in-vacuum case. Moreover, at a fixed $s$-wave interaction, with the increase of $p$-wave interaction, the lowest-lying phase changes from the three-body states with aabaab configuration to those with abbabb configuration. In other words, the in-vacuum three-body states with abbabb configuration become dominant in the regime with stronger $p$-wave interaction. It is because the increase of $p$-wave attraction induces larger enhancement for $bb$ pairings with smaller mass compared to $aa$ ones.
III.2 In-medium case
Figure 4: Contour plot of in-medium three-body energy $E_{3}^{aab}/E_{0}$ in a plane of mass $m_{a}/m_{0}$ and $m_{b}/m_{0}$. The mass $m_{0}$ and energy $E_{0}=k_{\rm F}^{2}/(2m_{0})$ are taken as the reference scales. The interaction strengths are fixed at $1/(k_{\rm F}a_{s})=1/(k_{\rm F}a_{p})=1.5$. The momentum cutoff is taken as $10k_{\rm F}$.
To see effects from the mass imbalance to the in-medium three-body ground-state energy, in Fig. 4, we plot the contour plot of in-medium three-body energy $E_{3}^{aab}/E_{0}$ in a plane of mass $m_{a}/m_{0}$ and $m_{b}/m_{0}$. For comparison, the mass $m_{0}$ and energy $E_{0}=k_{\rm F}^{2}/(2m_{0})$ are taken as the reference scales. The interaction strengths are fixed at $1/(k_{\rm F}a_{s})=1/(k_{\rm F}a_{p})=1.5$. The momentum cutoff $\Lambda$ is taken as $10k_{\rm F}$. In the present two-component Fermi gas, as for one species of atom with a fixed mass, if we gradually enlarge the mass of another one, the absolute value of $E_{3}^{aab}$ monotonically decreases. Such a qualitative behavior is a result of fixed interaction strengths. Moreover, it is also seen that a smaller reduced mass $m_{r}=m_{a}m_{b}/(m_{a}+m_{b})$ gives an negatively larger $E_{3}^{aab}$ in Fig. 4.
Figure 5: Lower-lying excited state of Cooper triple phases with different configurations in a plane of two-body $p$-wave interaction strengths. The mass ratio is taken as $m_{a}=2m_{b}$, and the momentum cutoffs are taken as $\Lambda/k_{\rm F}=10$.
In order to see the competition between the in-medium three-body correlations with different configurations in the present system, by fixing the mass ratio as $m_{a}=2m_{b}$, we plot the lower-lying excited state of Cooper triple phases with different configurations in a plane of two-body $p$-wave interaction strengths in Fig. 5. Here the dimensionless $p$-wave coupling strengths are defined as $\tilde{U}_{i}=2m_{r}k_{\rm F}U_{i}$. In this paper, we consider the attractive two-body $p$-wave interactions, which indicates that a larger absolute value of $\tilde{U}_{i}$ represents a stronger attraction. At a fixed $p$-wave interaction strength $\tilde{U}_{a}$ among atoms $a$, when the $p$-wave interaction strength $\tilde{U}_{b}$ among atoms $b$ increases to a certain value, the dominant (lower-lying) Cooper triple state changes from the aabaab configuration to the abbabb one. The similar behavior holds for a fixed $p$-wave interaction strength $\tilde{U}_{b}$ among atoms $b$. The competition between two different configurations of Cooper triples is similar to the in-vacuum case shown in Fig. 3, which results from the mass difference.
Figure 6: Phase diagram of $s$-wave pairing ($|E_{2,s}|>|E_{2,p}^{ii}|,\ |E_{3}^{iij}|$), $p$-wave pairings ($|E_{2,p}^{ii}|>|E_{2,s}|,\ |E_{3}^{iij}|$), and Cooper tripling ($|E_{3}^{iij}|>|E_{2,s}|,\ |E_{2,p}^{ii}|$) in the plane of $1/(k_{\rm F}a_{s})$ and $1/(k_{\rm F}a_{p})$ ($i,j=a,b$). The mass ratio is taken as $m_{a}=2m_{b}$, and the momentum cutoffs are taken as $\Lambda/k_{\rm F}=10$.
At last, by taking the mass ratio as $m_{a}=2m_{b}$, we summarize the ground-state phase diagram of $s$- and $p$-wave Cooper pairing states, and the Cooper triple states with different configurations in Fig. 6. The boundary lines between different kinds of Cooper clustering phases are determined in such a way that the one between tripling (iijiij configuration) and $s$-wave pairing, that between tripling (iijiij configuration) and $p$-wave pairing ($ii$ configuration), and that between $s$- and $p$-wave pairing ($ii$ configuration) are given by $E_{3}^{iij}=E_{2,s}$, $E_{3}^{iij}=E_{2,p}^{ii}$, and $E_{2,s}=E_{2,p}^{ii}$, respectively. The range of scattering parameters in Fig. 6 $0.5\leq 1/(k_{\rm F}a_{s})\leq 3$ and $1.5\leq 1/(k_{\rm F}a_{p})\leq 1.5$] are adopted as same one in Ref. [5], which are also similar values were realized in a recent experimental work [47]. Such a phase diagram captures interesting features associated with competing $s$- and $p$-wave pairings and moreover tripling. In order to investigate the physics behind more, we also plot in-medium clustering energies as functions of scattering lengths in Fig. 2. From Fig. 2, it is shown that the presence of the Fermi sea reshapes the formation of pairing and three-body states. Moreover, by comparing Fig. 2(a) [or (b)] and (c), the presence of Fermi surface affects three-body energies much more than pairing ones. Consequently, it is seen that when $s$- or $p$-wave interaction strength is comparably weak, the pairing phases survive, which is different to the in-vacuum case. As for the pairing states, the $p$-wave pair ($aa$ configuration) phase is always suppressed due to the mismatch of Fermi energies between atoms $a$ and $b$, where the Cooper instability among heavier atoms becomes weaker. As for the tripling states, the Cooper triple (aabaab configuration) phase survives due to the coexistence of both $s$- and $p$-wave interactions, where the effect from the mismatch of Fermi energies is partially offset by the inclusion of $s$-wave attraction. In the meanwhile, similar to the in-vacuum case shown in Fig. 3, there is a competition between Cooper triple (aabaab configuration) phase and Cooper triple (abbabb configuration) phase. On the one hand, we have numerically checked that the qualitative behaviors of the boundary line between $s$-wave pair phase and Cooper triple (aabaab configuration) phase, and that between Cooper triple (aabaab configuration) phase and Cooper triple (abbabb configuration) phase are unchanged even up to a extremely large $s$-wave interaction strength [such as $1/(k_{\rm F}a_{s})=10.0$]. It indicates that the competition between different configurations of Cooper triple phases may also result from the difference of scatterings between $s$-wave dimer ($ab$ configuration) and each component of atoms. In the meanwhile, at a fixed $s$-wave interaction strength, $bb$ configuration with a smaller mass also becomes more tightly bound than the $aa$ configuration as the $p$-wave scattering length increases due to the mismatch of Fermi energy between atoms $a$ and $b$. In other words, the Cooper triple (abbabb configuration) is expected to be stabler than the Cooper triple (aabaab configuration) in the stronger $p$-wave interaction regime. As for the cutoff dependence, we have numerically checked that while the regime of Cooper triple phases gets enlarged with larger momentum cutoff $\Lambda$, the boundaries between different kinds of pairings remain robust. Such a cutoff dependenc is also similar to that in the previous work [5]. Even though the specific value of momentum cutoff in the practical calculations is needed to compare our results with the experiments, our phase diagram would be useful to understand the qualitative features of hybridized $s$- and $p$-wave interacting systems with mass imbalance and different configurations of clustering.
IV Summary and perspectives
As a step further of previous work [5], in this paper we have investigated the mass imbalance and different kinds of clustering configurations in the system with the coexistence $s$- and $p$-wave interactions. We have investigated both the in-vacuum and in-medium two- and three-body correlations in the mass-imbalanced two-component Fermi gas. The solutions of in total five kinds of stable in-medium two- and three-body cluster states have been found in this system by solving the variational equations. By taking a certain mass ratio between two components of fermions, we have featured out a ground-state phase diagram consisting of $s$- and $p$-wave pairing states, and the Cooper triple states with different configurations in a plane of $s$- and $p$-wave scattering lengths.
With the inclusion of mass imbalance and different clustering configurations, our work would be useful for the deeper understanding and better understanding of the nontrivial states arising in the systems with the coexistence of both even-parity ($s$-wave) and odd-parity ($p$-wave) interactions. Present results would also be useful for the further investigation of unconventional superfluids and superconductors. For the future perspectives, it is also interesting to see how similar in-medium correlations may arise in other systems, such as condensed-matter or nuclear ones. For instance, our present model setup may be applied to the investigations of hypernuclei with the lightest hyperon $\Lambda$ [16].
Acknowledgements.
The author thanks Hiroyuki Tajima for useful discussions. Y.G. was supported by RIKEN Special Postdoctoral Researchers Program.